### v3

# The Effect of Omitted Variables on

Abstract
We show that, depending on how the impact of omitted variables is measured, it can be substantially easier for omitted variables to flip coefficient signs than to drive them to zero. This behavior occurs with “Oster’s delta” (Oster 2019), a widely reported robustness measure. Consequently, any time this measure is large—suggesting that omitted variables may be unimportant—a much smaller value reverses the sign of the parameter of interest. We propose a modified measure of robustness to address this concern. We illustrate our results in four empirical applications and two meta-analyses. We implement our methods in the companion Stata module regsensitivity.
JEL classification: C14; C18; C21; C25; C51
Keywords: Identification, Treatment Effects, Partial Identification, Sensitivity Analysis, Unconfoundedness
1 Introduction
The analysis of causality often relies on untestable assumptions, like the absence of unobserved omitted variables that could bias one’s findings. A large literature in statistics and econometrics now provides many tools that allow researchers to perform sensitivity analyses to assess the importance of these assumptions. Many of these methods use breakdown points as quantitative measures of the robustness of one’s conclusions to departures from the baseline identifying assumptions. These points tell us how much we can depart from our identifying assumptions before our empirical conclusions break down in some way.
In this paper, we distinguish between two specific kinds of breakdown points. The first is the explain away breakdown point. It answers the question
  • What is the smallest value of the sensitivity parameter required for the data to be consistent with a zero causal effect?
The second is the sign change breakdown point. It answers the question
  • What is the smallest value of the sensitivity parameter required for the data to be consistent with a causal effect that has a different sign from the causal effect we found in the baseline model?
These two breakdown points are often equal, but are not generally equivalent. In particular, the sign change breakdown point can be smaller than the explain away breakdown point. Hence it can be easier to reverse the sign of one’s results than to drive them to zero. This can occur when the omitted variable bias is discontinuous in the sensitivity parameter, allowing the value of the bias adjusted estimand to jump across the horizontal axis at zero as the sensitivity parameter varies. This behavior is illustrated in figure 1, which we discuss more below. Such discontinuities can arise in regression analysis because the sensitivity parameters often involve covariance and variance terms, which lead to nonlinear restrictions on the value of the bias.
In this paper, we study the relationship between these two kinds of breakdown points in coefficient stability analyses used to assess the importance of omitted variables in linear regression analysis. These analyses examine how coefficients change when additional regressors are added, and have a long history in empirical economics. Folk wisdom holds that if the additional covariates increase the R-squared substantially, but the coefficient nonetheless does not change much, then we can be more confident that any further omitted variable bias is small. For example, see the textbook discussion on pages 74–78 of Angrist and Pischke (2015). This idea was first formalized by Altonji et al. (2005a, b, 2008). It was later substantially extended by Oster (2019) (hereafter Oster), an extremely influential paper with about 4500 Google Scholar citations as of December 2024. Finkelstein et al. (2021, AER, page 2706) describe it as providing “the now-standard methodology” and “the standard approach” to adjusting for selection on unobservables. [^3] Other papers on regression sensitivity include Mauro (1990), Murphy and Topel (1990), Frank (2000), Imbens (2003), Altonji et al. (2005a), Clarke (2009), Bellows and Miguel (2009), Hosman et al. (2010), Gonzalez and Miguel (2015), Krauth (2016), Cinelli and Hazlett (2020), and Diegert et al. (2022) for linear models and Rosenbaum and Rubin (1983), Rosenbaum (1995, 2002), Robins et al. (2000), Altonji, Elder, and Taber (2005a, 2008), Masten and Poirier (2018, 2020), and Altonji et al. (2019) for nonlinear models. Moreover, from 2019–2021, 23 papers published in the top five economics journals formally compute and report the results from this method, while another 8 papers informally reference this method in support of their analysis. Based on the data collected by Blandhol et al. (2022), this is more than the number of papers that use 2SLS per year, among papers published in top five economics journals. We discuss in detail how Oster’s (2019) results are used in empirical economics in appendix LABEL:sec:inPractice.
In section 2 we review Oster’s (2019) method. The main sensitivity parameter for this method, denoted by $\delta$, is commonly interpreted as the ratio of the magnitude of selection on unobservables to the magnitude of selection on observables. We then study breakdown points for this sensitivity parameter in section 3. Oster’s Proposition 2 characterizes the explain away breakdown point. Following her recommendation, empirical researchers now routinely report estimates of this explain away breakdown point as a measure of robustness to omitted variables. We show that this explain away breakdown point is very different from the sign change breakdown point. In fact, we show that any time the explain away breakdown point is large—suggesting that omitted variables may be unimportant—a much smaller value of her sensitivity parameter can actually change the sign of the parameter of interest. In particular, we prove that the sign change breakdown point is bounded above by $\delta=1$. This is a substantial concern since 1 is often viewed as the cutoff for a robust result. Consequently, if we maintain 1 as the cutoff, our result implies that no empirical results are robust to sign changes, using Oster’s method. Put differently: The true causal effect can have the opposite sign as the baseline estimand any time the magnitude of selection on unobservables is at least as large as the magnitude of selection on observables.
Figure 1: Example of Oster’s identified set $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})$ for the regression coefficient $\beta$ as a function of the sensitivity parameter $\delta$ (see section 2 for a definition of this set and a description of the $R_{\text{long}}^{2}$ parameter). As we discuss in section 2.3, this identified set always has either one, two, or three elements. Both plots show the same set, but with different horizontal axis ranges. In this example, the explain away breakdown point is $|-32|$. This is shown in the magnified region on the left plot. In contrast, the sign change breakdown point is 0.586, as shown in the right plot. The plot also shows the vertical asymptote at $\delta=1$ which we discuss in section 3.1.
To build intuition for our main results in section 3, consider figure 1. Both plots show the same curve, but with different horizontal axis ranges, to highlight different aspects which we discuss below. This curve is an example from our empirical application in section 6. It shows the pairs of regression coefficient values $\beta$ (on the vertical axis) and Oster’s main sensitivity parameter $\delta$ (on the horizontal axis) that are consistent with the data and assumptions. Importantly, as shown in the right plot, for some values of $\delta$ there are multiple values of $\beta$ that are consistent with the data. This occurs because $\delta$ is defined to be a ratio of two regression coefficients, and hence knowledge of $\delta$’s exact value implies that $\beta$ must satisfy a fairly complicated nonlinear constraint. Specifically, it must be the solution to a cubic equation (see equation (4) on page 4), which can have multiple roots. This explains the source of the discontinuity in the value of the omitted variable bias that we discussed earlier, and why for this sensitivity parameter it is easier to flip the coefficient’s sign than to obtain an exact zero.
Concretely, in this example the baseline estimate is on the vertical intercept at $\delta=0$. This estimate is positive. How robust is this estimate to the presence of omitted variables? In the left plot we see that we must go all the way to $\delta=-32$ before the value $\beta=0$ is consistent with the data and assumptions. This is the number that researchers commonly report as a measure of robustness. However, from the right plot, we see that we only need to go to $\delta=0.586$ to find a value of $\beta$ that is negative. This value $\delta=0.586$ is the sign change breakdown point. Oster (2019) does not discuss this breakdown point, its companion Stata package psacalc does not compute it, and none of the top five papers we survey in appendix LABEL:sec:inPractice report it. Yet, as we see here, the sign change breakdown point is substantially smaller than $|-32|$, the explain away breakdown point. Indeed, it is below $\delta=1$, the commonly used cutoff for a robust result. Consequently, in this empirical example, focusing on the sign change breakdown point overturns the authors’ conclusion of robustness: Only a small amount of selection on unobservables relative to observables is necessary to flip the sign of the baseline estimate.
Importantly, all of the coefficient values plotted in figure 1, even the seemingly extreme values, are equally consistent with the data and assumptions. Under the assumptions of Oster’s Proposition 2, which is used to obtain these plots, there is no justification for ignoring any of these values. The only way to eliminate values is by making additional assumptions. One reasonable assumption, which we consider in section 3.2, is to restrict the magnitude of the omitted variable bias. With this additional assumption, we can justifiably eliminate some of the most extreme values. This assumption does not overturn our main conclusions, however. For example, in figure 1 the coefficient value consistent with $\delta=0.586$ is not extreme at all—it is the same order of magnitude as the baseline estimate. So in this example, the sign change breakdown point is still very different from the explain away breakdown point, even with a substantial restriction on the magnitude of the omitted variable bias. Moreover, magnitude restrictions require researchers to choose a tuning parameter—just how much should they restrict the magnitude? If the magnitude is restricted too much, then this assumption is essentially equivalent to assuming a priori that there is no omitted variable bias, in which case a sensitivity analysis would be unnecessary. But as soon as reasonably sized magnitudes are allowed, the concerns we discuss in sections 3 and 4 arise.
Refer to caption         Refer to caption
Figure 2: Left: The cumulative identified set for the example in figure 1. Here we see that for large values of $\bar{\delta}$ the identified set includes every value in $\mathbb{R}$ except a small interval around 0. The set includes 0 for $\bar{\delta}\geq 32$, the explain away breakdown point. But the sign change breakdown point occurs much earlier, at $\delta=0.586$. Moreover, note that the identified set at $\bar{\delta}=1$ is $(-\infty,-0.0855]\cup[0.0432,1.8947]$. At $\bar{\delta}=1.05$ it is $(-\infty,-0.0798]\cup[0.0415,\infty)$. Right: The convex hull of the left plot.
Researchers also commonly report selection bias adjusted estimates. In section 4 we show that the most commonly used adjustment is extremely sensitive to perturbations of Oster’s sensitivity parameter. To see why, consider again figure 1. There is always an asymptote at $\delta=1$. This implies that bias adjustments computed at $\delta$ exactly equal to one—which is currently standard empirical practice—are not representative of bias adjustments computed at nearby values of $\delta$.
To address the issues described in sections 3 and 4, we propose several simple modifications to the methods in Oster (2019). First, we recommend gathering all values of the coefficient consistent with a range of sensitivity parameters, and presenting this set. The left plot in figure 2 gives an example: Here we gather all $\beta$’s consistent with $|\delta|$ less than some fixed $\bar{\delta}$. We then plot that set as a function of $\bar{\delta}$. The graph of these values of $\beta$ is discontinuous in $\bar{\delta}$ at $\bar{\delta}=0.586$, which explains the difference in the sign change breakdown and explain away breakdown points. As we saw in figure 1, nearly all values of $\beta$, including arbitrarily large negative values, are consistent with an upper bound of $\bar{\delta}=1.05$, but a small range around 0 is not part of this set until $\bar{\delta}$ equals 32. Researchers who want to focus only on the largest and smallest possible values of $\beta$ can instead present the convex hull, which is shown in the right plot of figure 2.
Second, we develop the sign change breakdown point for Oster’s main sensitivity parameter and recommend that researchers report this rather than, or in addition to, the explain away breakdown point. Furthermore, in section 3.2, we provide extensions to these two recommendations to formally incorporate the magnitude restrictions discussed above. We show that the sign change breakdown point can be strictly larger than 1 when a magnitude restriction is imposed. Therefore it is possible to recover findings of robustness even without changing the cutoff for robustness. This is not always the case however, as we saw in figure 1.
In section 6 we illustrate our results using data from Satyanath et al. (2017, Journal of Political Economy), who study the effect of social capital on the rise of the Nazi party. There we show just how different the sign change and explain away breakdown points can be in practice. We then show how magnitude restrictions can sometimes help retain findings of robustness. We also illustrate the extreme sensitivity of bias adjusted estimates to perturbations of the sensitivity parameter. In appendix LABEL:sec:extraEmpirical we study three additional empirical applications in detail, all published in the American Economic Review.
To see that these four applications are not outliers, in section 7 we conduct two meta-analyses covering 58 different empirical applications published in the top five economics journals plus the AEJ: Applied Economics. We show that the distinction between explain away and sign change breakdown points leads to substantively different conclusions about robustness to omitted variables in the majority of empirical applications. In particular, suppose researchers were to continue the current standard practice of reporting explain away breakdown points, but wanted to interpret them as sign change breakdown points. We show that in more than 50% of applications, this interpretation is only valid if researchers already know the sign of coefficient of interest. But in that case, it would not be necessary to do a sensitivity analysis, since the result is known a priori.
Thus far we have discussed ways of assessing omitted variable bias by using $\delta$ as the main sensitivity parameter. We also briefly discuss several critiques of $\delta$ itself in section 5. Finally, in section 8 we conclude with a summary of our recommendations.
2 Setup and Identification Analysis
In this section we review the main identification results in Oster (2019), including her breakdown analysis. Throughout this section, as well as sections 35, we abstract from sampling uncertainty and assume the population data is known. We briefly discuss estimation and inference in appendix LABEL:sec:estimation.
2.1 Regressions of Interest
Let $Y$ be an outcome variable, $X$ a treatment variable, $W_{1}$ a vector of observed covariates, and $W_{2}$ a vector of unobserved covariates. The following assumption ensures that the OLS estimands we will consider are well defined.
Assumption A1.
$\operatorname*{var}(Y,X,W_{1},W_{2})$ is finite. $\operatorname*{var}(X,W_{1},W_{2})$ and $\operatorname*{var}(Y,X,W_{1})$ are positive definite.
A1 allows $Y$ to be perfectly collinear with $(X,W_{1},W_{2})$, but not with $(X,W_{1})$. Collinearity of $Y$ with $(X,W_{1})$ would violate A3 below.
Consider three OLS regressions:
  1. 1.
    The short regression of
    $Y\text{ on }(1,X).$
    Let $\beta_{\text{short}}$ denote the corresponding population regression coefficient on $X$. Let $R_{\text{short}}^{2}$ denote the corresponding R-squared.
  2. 2.
    The medium regression of
    $Y\text{ on }(1,X,W_{1}).$
    Let $\beta_{\text{med}}$ denote the corresponding population regression coefficient on $X$. Let $R_{\text{med}}^{2}$ denote the corresponding R-squared. Let $\gamma_{1,\text{med}}$ denote the coefficients on $W_{1}$.
  3. 3.
    The long regression of
    $Y\text{ on }(1,X,W_{1},W_{2}).$
    Let $\beta_{\text{long}}$ denote the corresponding population regression coefficient on $X$. Let $R_{\text{long}}^{2}$ denote the corresponding R-squared. Note that $R_{\text{long}}^{2}\in[R_{\text{med}}^{2},1]$. Let $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$ denote the coefficients on $W\coloneqq(W_{1},W_{2})$.
Suppose $\beta_{\text{long}}$ is the parameter of interest. In particular, we can write the long regression as
$Y=\beta_{\text{long}}X+\gamma_{1,\text{long}}^{\prime}W_{1}+\gamma_{2,\text{long}}^{\prime}W_{2}+Y^{\perp X,W}$ (1)
where $Y^{\perp X,W}$ is defined to be the OLS residual plus the intercept term, and hence is uncorrelated with each component of $(X,W_{1},W_{2})$ by construction. The identification problem is that $W_{2}$ is not observed. Instead, we only assume that the joint distribution of $(Y,X,W_{1})$ is known. This allows us to point identify $\beta_{\text{short}}$ and $\beta_{\text{med}}$. These coefficients usually do not equal $\beta_{\text{long}}$, however. The difference between $\beta_{\text{long}}$ and either $\beta_{\text{short}}$ or $\beta_{\text{med}}$ is omitted variable bias. We next consider assumptions that will help us bound the magnitude of this omitted variable bias.
2.2 Comparing Selection on Observables and on Unobservables
Following the analysis of Altonji et al. (2005a, pages 175–176), Oster (2019) recommends that we measure the magnitude of selection on unobservables via the parameter
$\delta\coloneqq\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}^{\prime}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}^{\prime}W_{2})}\Bigg{/}\frac{\operatorname*{cov}(X,\gamma_{1,\text{long}}^{\prime}W_{1})}{\operatorname*{var}(\gamma_{1,\text{long}}^{\prime}W_{1})}.$ (2)
To allow for cases where the denominator is zero, our formal analysis uses the following assumption.
Assumption A2.
There is a known $\delta\in\mathbb{R}$ such that
$\delta\cdot\frac{\operatorname*{cov}(X,\gamma_{1,\text{long}}^{\prime}W_{1})}{\operatorname*{var}(\gamma_{1,\text{long}}^{\prime}W_{1})}=\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}^{\prime}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}^{\prime}W_{2})}.$ (3)
The following assumption, combined with A1, ensures that we are not dividing by zero in equation (3).
Assumption A3.
$\gamma_{1,\text{long}}\neq 0$ and $\gamma_{2,\text{long}}\neq 0$.
The parameter $\delta$ depends on two terms:
  1. 1.
    (“Selection on unobservables”) Regress $X$ on $(1,\gamma_{2,\text{long}}^{\prime}W_{2})$ and get the coefficient on the index $\gamma_{2,\text{long}}^{\prime}W_{2}$.
  2. 2.
    (“Selection on observables”) Regress $X$ on $(1,\gamma_{1,\text{long}}^{\prime}W_{1})$ and get the coefficient on the index $\gamma_{1,\text{long}}^{\prime}W_{1}$.
$\delta$ is the ratio of these two regression coefficients. $\delta$ is not known from the data since it depends on $W_{2}$, which is not observed, and $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$, which are also unknown. Instead, we will study what can be said about $\beta_{\text{long}}$ under assumptions on $\delta$. Several papers propose alternative sensitivity parameters, including Cinelli and Hazlett (2020) and Diegert et al. (2022). In this paper we instead use the same sensitivity parameter $\delta$ as Oster (2019), since our focus here is to study the implications of using different kinds of breakdown points. We give more discussion of $\delta$’s interpretation in section 5 below.
Baseline covariates
$\delta$ is defined to calibrate the magnitude of selection on unobservables by using the observed covariates $W_{1}$. Typically we will have additional covariates $W_{0}$ which we want to include in the analysis, but not use for calibration. Denote these baseline controls by $W_{0}$. To incorporate these covariates, replace $(X,W_{1},W_{2})$ with $(X^{\perp W_{0}},W_{1}^{\perp W_{0}},W_{2}^{\perp W_{0}})$ throughout the entire analysis. See section 3.3 of Diegert et al. (2022) for a justification of this approach and further discussion. For simplicity, we omit $W_{0}$ throughout our theoretical analysis.
2.3 Identification of $\beta_{\text{long}}$
In this section we state a version of the main and most general identification result in Oster (2019). That paper also analyzes several special cases which we do not discuss for brevity. The main substantive assumption is the following.
Assumption A4 (Exogenous controls).
$\operatorname*{cov}(W_{1},W_{2})=0$.
A4 requires that all of the observed covariates are uncorrelated with the omitted variable. Appendix A.1 in Oster (2019) briefly discusses one approach to relaxing A4, which we discuss in section 5. Also see Appendix A of Diegert et al. (2022) for additional details. Diegert et al. (2022) also develop an alternative approach to sensitivity analysis that does not rely on A4. In this paper we maintain A4 throughout because our main focus is to study different notions of breakdown. For simplicity, from here on we assume $W_{2}$ is a scalar. All of our results can be generalized to vector $W_{2}$ using the arguments in Appendix B of Diegert et al. (2022).
To characterize the possible values for $\beta_{\text{long}}$, it is useful to consider the regression of $X$ on $(1,W_{1})$. Let $\pi_{1}$ denote the coefficient on $W_{1}$ and let $X^{\perp W_{1}}\coloneqq X-\pi_{1}^{\prime}W_{1}$.
Under the above assumptions, Oster’s Proposition 2 shows that $\beta_{\text{long}}$ must lie in the set
$\mathcal{B}(\delta,R_{\text{long}}^{2})\coloneqq\{b\in\mathbb{R}:f(\beta_{\text{med}}-b,\delta,R_{\text{long}}^{2})=0\},$ (4)
where $f(B,\delta,R_{\text{long}}^{2})\coloneqq f_{0}(B)+\delta\cdot f_{1}(B,R_{\text{long}}^{2})$ with
$f_{0}(B)$ $\coloneqq-B\operatorname*{var}(X^{\perp W_{1}})\left(\operatorname*{var}(\gamma_{1,\text{med}}^{\prime}W_{1})+2B\operatorname*{cov}(X,\gamma_{1,\text{med}}^{\prime}W_{1})+B^{2}\operatorname*{var}(\pi_{1}^{\prime}W_{1})\right)$
$f_{1}(B,R_{\text{long}}^{2})$ $\coloneqq(R_{\text{long}}^{2}-R_{\text{med}}^{2})\operatorname*{var}(Y)\operatorname*{cov}(X,\gamma_{1,\text{med}}^{\prime}W_{1})+B(R_{\text{long}}^{2}-R_{\text{med}}^{2})\operatorname*{var}(Y)\operatorname*{var}(\pi_{1}^{\prime}W_{1})$
$\qquad+B^{2}\operatorname*{var}(X^{\perp W_{1}})\operatorname*{cov}(X,\gamma_{1,\text{med}}^{\prime}W_{1})+B^{3}\operatorname*{var}(X^{\perp W_{1}})\operatorname*{var}(\pi_{1}^{\prime}W_{1}).$
As Oster (2019) discusses, the set in (4) are the roots of a cubic polynomial. Hence this set has at most three elements. Figure 1 shows an example of this set, as a function of $\delta$. Our Theorem LABEL:thm:IDset in Appendix LABEL:sec:technical extends this analysis to show that $\mathcal{B}(\delta,R_{\text{long}}^{2})$ is essentially sharp. That is, almost any element in this set is in fact a feasible value of $\beta_{\text{long}}$. Formally, we show that it equals the identified set (denoted by $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})$), after subtracting off the set $\mathcal{B}_{\text{A3fail}}\coloneqq\{b\in\mathbb{R}:\gamma_{1,\text{med}}+(\beta_{\text{med}}-b)\pi_{1}=0\}$:
$\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})=\mathcal{B}(\delta,R_{\text{long}}^{2})\setminus\mathcal{B}_{\text{A3fail}}.$ (5)
The set we subtract off are values of $\beta_{\text{long}}$ that are not consistent with A3; this set is typically empty, however. This sharpness is important to ensure that breakdown analysis is not unnecessarily conservative. Finally, in Appendix LABEL:sec:technical we also show how the analysis simplifies to the baseline case when $\delta=0$, which gives $\beta_{\text{long}}=\beta_{\text{med}}$.
2.4 Oster’s Breakdown Analysis
Thus far we have characterized the identified set for $\beta_{\text{long}}$ when $\delta$ and $R_{\text{long}}^{2}$ are known. Next we consider breakdown analysis. In this section we review the main breakdown point result in Oster (2019). We then compare it with the sign change breakdown point in section 3.
Let $\beta_{\text{hypo}}\in\mathbb{R}$. Define the exact value breakdown point
$\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})\coloneqq\inf\{|\delta|:\delta\in\mathbb{R},\beta_{\text{hypo}}\in\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\}.$
This is the smallest magnitude of the sensitivity parameter $\delta$ that is compatible with $\beta_{\text{long}}=\beta_{\text{hypo}}$. That is, for all $\delta$ with $|\delta|<\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})$, we know that $\beta_{\text{hypo}}$ is not in the identified set $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})$. The next result gives a closed form solution for this value.
Theorem 1.
Suppose the joint distribution of $(Y,X,W_{1})$ is known, A1, A3, and A4 hold, $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known, $\beta_{\text{long}}=\beta_{\text{hypo}}$ for a known $\beta_{\text{hypo}}\in\mathbb{R}\setminus\mathcal{B}_{\text{A3fail}}$, and $\beta_{\text{short}}\neq\beta_{\text{med}}$. Then the solution $\delta$ to equation (3) is point identified and equals
$\frac{-f_{0}(\beta_{\text{med}}-\beta_{\text{hypo}})}{f_{1}(\beta_{\text{med}}-\beta_{\text{hypo}},R_{\text{long}}^{2})}.$
Consequently, the absolute value of this term equals $\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})$.
Theorem 1 is very similar to Oster’s Proposition 3, with only minor differences. As Oster remarked, this result shows that, even though knowledge of $\delta$ does not point identify $\beta_{\text{long}}$, knowledge of $\beta_{\text{long}}$ does point identify $\delta$. Figure 1 illustrates why: Fixing $\delta$ is equivalent to placing a vertical line on the plot. This vertical line may intersect the displayed curve up to three times, since $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})$ can have up to three elements. Fixing the value $\beta_{\text{long}}$ to be equal to some known $\beta_{\text{hypo}}$, however, is equivalent to placing a horizontal line on the plot. This line can only intersect the displayed curve once. The absolute value of the horizontal coordinate at this intersection is precisely $\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})$.
Setting $\beta_{\text{hypo}}=0$, we get the explain away breakdown point
$\delta^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})\coloneqq\delta^{\text{bp},=}(0,R_{\text{long}}^{2})=\frac{|f_{0}(\beta_{\text{med}})|}{|f_{1}(\beta_{\text{med}},R_{\text{long}}^{2})|}.$
Its signed value, $-f_{0}(\beta_{\text{med}})/f_{1}(\beta_{\text{med}},R_{\text{long}}^{2})$, is often called Oster’s delta. This is the unique value of $\delta$ that is compatible with an exactly zero long regression coefficient, $\beta_{\text{long}}=0$. Estimates of this value are commonly reported as a measure of the robustness of the baseline model $(\delta=0)$ to the presence of omitted variables.
3 Measuring the Robustness of $\text{sign}(\beta_{\text{long}})$ to Omitted Variables
In this section we analyze the robustness of conclusions about the sign of $\beta_{\text{long}}$ to the presence of omitted variables. In section 3.1 we first show that the sign change breakdown point for $\delta$ can never be larger than 1. This is a concern since $|\delta|=1$ is often considered the cutoff for robustness, and therefore this result implies that conclusions about the sign of $\beta_{\text{long}}$ are always non-robust. In section 3.2 we propose a simple modification to Oster’s analysis that allows the sign change breakdown point to be larger than 1.
3.1 The Sign Change Breakdown Point for $\delta$
Oster’s delta answers the following question:
  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $\delta$ that is required for the data to be consistent with $\beta_{\text{long}}=0$?
We call this the explain away breakdown point. Researchers may also be interested in answering the following question:
  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $\delta$ such that $\beta_{\text{long}}$ has a different sign from $\beta_{\text{med}}$?
We call this the sign change breakdown point. As mentioned earlier, the answers to these two questions can be different. In fact, as we next show, the sign change breakdown point is typically much smaller than the explain away breakdown point.
To formalize this claim, define the sign change breakdown point for a given value of $R_{\text{long}}^{2}$ as
$\delta^{\text{bp},\text{sign}}(R_{\text{long}}^{2})\coloneqq\begin{cases}\delta^{\text{bp},>}(R_{\text{long}}^{2})&\text{if $\beta_{\text{med}}>0$}\\ \delta^{\text{bp},<}(R_{\text{long}}^{2})&\text{if $\beta_{\text{med}}<0$}\end{cases}$
where
$\delta^{\text{bp},>}(R_{\text{long}}^{2})$ $\coloneqq\inf\{|\delta|:\delta\in\mathbb{R},b\leq 0\text{ for some }b\in\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\}$
$\delta^{\text{bp},<}(R_{\text{long}}^{2})$ $\coloneqq\inf\{|\delta|:\delta\in\mathbb{R},b\geq 0\text{ for some }b\in\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\}.$
We now state our main theoretical result.
Theorem 2.
Suppose A1A4 hold. Suppose $\beta_{\text{short}}\neq\beta_{\text{med}}$. Then for any $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$,
$\delta^{\text{bp},\text{sign}}(R_{\text{long}}^{2})\leq 1.$
Here we assume $\beta_{\text{short}}\neq\beta_{\text{med}}$ for simplicity only. Theorem 2 shows that, in Oster’s sensitivity analysis, the sign change breakdown point can never be larger than one. In practice, researchers typically consider the value $\delta=1$ to be the cutoff for determining the robustness of a result. That is, they interpret values of $\delta$ greater than 1 as robust and values less than 1 as sensitive. Based on this criterion, Theorem 2 therefore shows that the sign of the parameter of interest is never robust (assuming we require $\delta$ strictly larger than 1 to conclude that the result is robust). In contrast, the explain away breakdown point $\delta^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ is often substantially larger than one. We illustrate this difference empirically in section 6.
Figure 1 illustrates this result: The graph $\{(\delta,b)\in\mathbb{R}^{2}:f(\beta_{\text{med}}-b,\delta,R_{\text{long}}^{2})=0\}$ has an asymptote at $\delta=1$. Consequently, values of $\delta$ close to one—but not equal to one—yield identified sets that contain arbitrarily large or arbitrarily small values. Hence for $\delta$ values near one we cannot point identify the sign of $\beta_{\text{long}}$. However, much larger values of $\delta$ may be needed before the graph crosses the horizontal axis $b=0$ exactly. This explains why the explain away breakdown point can often be much larger than the sign change breakdown point for this model.
Intuition
The asymptote at $\delta=1$ arises because the sensitivity analysis allows the treatment $X$ and the covariates $(W_{1},W_{2})$ to be nearly multi-collinear. And, as we explain here, this near multi-collinearity implies that the omitted variable bias will be large while at the same time $\delta$ will be close to 1. We prove this formally in Proposition LABEL:prop:formal_asym_intuition in Appendix LABEL:sec:intuitionAppendix. To see it intuitively, consider a distribution of $(Y,X,W_{1},W_{2})$ that is consistent with the identified covariance matrix $\operatorname*{var}(Y,X_{1},W_{1})$ but where the omitted variable $W_{2}$ almost fully accounts for the variation in treatment after adjusting for the observed variables $W_{1}$. There are three main steps:
  1. 1.
    In this case, $X$, $W_{1}$, and $W_{2}$ are close to multi-collinear. Consequently, the coefficients in the long regression, $(\beta_{\text{long}},\gamma_{1,\text{long}},\gamma_{2,\text{long}})$, will be large. We prove that this is a general feature of near multi-collinear OLS estimands in Appendix LABEL:sec:OLSmultiCollinear. If $\beta_{\text{long}}$ is large then the OVB must be large because $\beta_{\text{med}}$ is fixed (that is, its value is not affected by the omitted variable $W_{2}$).
  2. 2.
    $\delta$’s measure of selection on unobservables is the coefficient from regressing $X$ on $\gamma_{2,\text{long}}W_{2}$:
    $\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2})}.$
    $\gamma_{2,\text{long}}$, as one of the long regression coefficients, will also be large because of near multi-collinearity. Consequently, this measure of selection on unobservables will be close to zero since regression coefficients scale down when the covariate is scaled up.
  3. 3.
    For the same reason, $\delta$’s measure of selection on observables will also be close to zero, since $\gamma_{1,\text{long}}$ will be large. Since the numerator and denominator of $\delta$ are both small, it is not obvious how the ratio will behave. However, it turns out that both terms converge to zero at the same rate, which implies that $\delta$ converges to a constant. This follows from the OVB formula for the coefficient on $W_{1}$:
    $\gamma_{1,\text{long}}-\gamma_{1,\text{med}}=(\beta_{\text{med}}-\beta_{\text{long}})C_{1}$
    where $C_{1}$ is a constant, and the OVB formula for the coefficient on $W_{2}$:
    $\gamma_{2,\text{long}}-0=(\beta_{\text{med}}-\beta_{\text{long}})C_{2}$
    where $C_{2}$ is a value that converges to a constant. We derive these equations in lemma LABEL:lemma:regfacts. Thus both $\gamma_{2,\text{long}}$ and $\gamma_{1,\text{long}}$ diverge at the same rate, controlled by the divergence of $\beta_{\text{long}}$. The divergence of these terms is what drives the numerator and denominator of $\delta$ to zero. Finally, some additional regression algebra shows that the constant limit of $\delta$ is one.
Thus in the first step we see that the OVB will be large while in the third step we see that $\delta$ will be close to one. This is precisely what happens along the asymptote at $\delta=1$. This followed from (a) analyzing the behavior of near multi-collinear OLS coefficients and (b) exploring the specific structure of $\delta$. Thus the asymptote arises because $(X,W_{1},W_{2})$ are allowed to be arbitrarily mutli-collinear. Researchers may be interested in explicitly ruling out this collinearity. This is exactly what the alternative methods of Cinelli and Hazlett (2020) and Diegert et al. (2022) do. Both papers include sensitivity parameters that explicitly bound the joint explanatory power of the observed and unobserved covariates in the first stage.
3.2 Incorporating Magnitude Restrictions on the Omitted Variable Bias
We have just shown that the sign change breakdown point for $\delta$ can never be larger than one. This negative result can be overturned by adding additional identifying assumptions. For example, there are usually values of the coefficient that are extremely implausible, such as a 500% return to education. Any procedure that produced such an estimate would therefore suggest that there is a problem with the procedure itself, the data, or both. In this section we show that if researchers are willing to explicitly impose this kind of bound then we can obtain sign change breakdown points larger than one. Specifically, consider the following assumption.
Assumption A5.
$|\beta_{\text{long}}-\beta_{\text{med}}|\leq M$ for some known $M\geq 0$.
A5 says that the magnitude of the omitted variable bias is no larger than $M$. This assumption can be viewed as formalizing Oster’s (2019) statement that researchers may be “willing to assume that the bias is fairly small” (page 194). The value $M$ should be chosen to reflect the researcher’s beliefs about the largest possible bias in their baseline estimand $\beta_{\text{med}}$. For example, consider a regression of log earnings $Y$ on years of schooling $X$ and other observed covariates $W_{1}$. Suppose our estimate of $\beta_{\text{med}}$ is 0.1. Then $M=1$ would say that the true coefficient on schooling—as obtained after adjusting for the unobserved confounders—is at most 1.1. This value is far outside the range of most estimates (e.g., Card 2001 table II) and so in this application researchers would likely view A5 with $M=1$ as a plausible assumption.
In general, a simple approach to selecting $M$ is to let it be a multiple of $\beta_{\text{med}}$. That is, let $M=p\cdot|\beta_{\text{med}}|$ for $p\geq 0$. For this choice, A5 is equivalent to
$\beta_{\text{long}}\in\big{[}\beta_{\text{med}}-p|\beta_{\text{med}}|,\beta_{\text{med}}+p|\beta_{\text{med}}|\big{]}.$
Every element of this set has the same sign if and only if $p<1$. For this reason, since we are interested in robustness to sign changes, $p$ should be chosen to be larger than 1. In the returns to schooling example with $\beta_{\text{med}}=0.1$, $M=1$ is equivalent to the choice $p=10$. We recommend that researchers consider a range of magnitude restrictions, as we illustrate empirically in section 6. For example, in section 6 we consider $p=2$, $3$, and $10$.
The next result shows the impact of this magnitude restriction on identification of $\beta_{\text{long}}$.
Corollary 1.
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1A5 hold. Suppose $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known. Recall that A2 specifies that $\delta\in\mathbb{R}$ is known. Then the identified set for $\beta_{\text{long}}$ is
$\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M)\coloneqq\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\cap[\beta_{\text{med}}-M,\beta_{\text{med}}+M].$
Let $\delta^{\text{bp},\text{sign}}(R_{\text{long}}^{2},M)$ denote the corresponding sign change breakdown point. For $0\leq M<\infty$, it is now possible for $\delta^{\text{bp},\text{sign}}(R_{\text{long}}^{2},M)>1$. We illustrate this empirically in section 6.
Thus far we have followed Oster (2019) by assuming the exact value of $\delta$, the measure of selection is known. We extend previous results by considering a known bound on the measure of selection, rather than its exact value. Specifically, we make the following assumption.
Assumption A6.
There is a known constant $\bar{\delta}\geq 0$ such that $|\delta|\leq\bar{\delta}$ for $\delta$ that satisfies equation (3).
This leads to the following identification result.
Corollary 2.
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1, A3, A4, and A6 hold. Suppose $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known. Then the identified set for $\beta_{\text{long}}$ is
¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})\coloneqq\bigcup_{|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2}).$
Suppose A5 also holds. Then the identified set for $\beta_{\text{long}}$ is
¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2},M)\coloneqq\bigcup_{|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M).$
This corollary simply notes that we take the union of the identified sets from equation (5) over all allowed values of the sensitivity parameter $\delta$ with magnitude less than $\bar{\delta}$ to obtain the identified set under A6.
4 The Sensitivity of Bias Adjusted Estimands
Thus far we have focused on breakdown analysis. Researchers also commonly compute point estimates that adjust for omitted variable bias. See our survey in appendix LABEL:sec:inPractice for examples from the empirical economics literature. The most commonly used bias adjustment is an estimate of
$\beta_{\text{long}}=\beta_{\text{med}}+(\beta_{\text{med}}-\beta_{\text{short}})\frac{R_{\text{long}}^{2}-R_{\text{med}}^{2}}{R_{\text{med}}^{2}-R_{\text{short}}^{2}}.$ (6)
Oster’s Proposition 1 shows that this equation holds if $\delta=1$ and an auxiliary assumption holds (see LABEL:assump:OsterA2 in Appendix LABEL:sec:BiasCorrectionsAppendix). However, equation (6) does not hold, even in an approximate sense, if $\delta$ is very close to one, but not exactly equal to one. Instead, $\beta_{\text{long}}$ can be arbitrarily large even for values of $\delta$ very close to one. As discussed earlier, this is due to the vertical asymptote in the identified set that occurs precisely at $\delta=1$. Consequently, bias adjustments based on equation (6) do not accurately represent the magnitude of omitted variable bias for values of $\delta$ close to one. We illustrate this discrepancy in table 6 in our empirical application. We discuss this issue in more detail in Appendix LABEL:sec:BiasCorrectionsAppendix. We also discuss several other issues with how bias adjustment is done in practice in Appendix LABEL:sec:commonPitfallsAppendix.
Rather than performing bias adjustments, researchers can plot the identified set ¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ for $\beta_{\text{long}}$ as a function of $\bar{\delta}\geq 0$. We described this set in section 3.2. This set describes the values of $\beta_{\text{long}}$ consistent with any amount of selection on unobservables $\delta$ with magnitude less than $\bar{\delta}$. Hence, unlike a single point estimand for $\delta=1$, it shows how the feasible values of $\beta_{\text{long}}$ vary for a range of $\delta$ values. See figure 2 for an example.
5 Interpreting $\delta$ and Additional Critiques
We have focused on the properties of various ways of assessing omitted variable bias that use $\delta$ as the main sensitivity parameter. The previous literature has also critiqued the use of $\delta$ itself as a sensitivity parameter. We briefly discuss those and several other critiques in this section.
5.1 $\delta$ is a Double Ratio
Cinelli and Hazlett (2020, section 6.3) show that $\delta$ can be interpreted as a double ratio, and use this result to argue that $\delta$ does not accurately measure selection on unobservables versus selection on observables. We summarize their analysis here. For simplicity, consider the case where $W_{1}$ and $W_{2}$ are scalar and are normalized to have variance 1. Maintain assumptions A1A4 from earlier. Let
$X$ $=\pi_{1}W_{1}+\pi_{2}W_{2}+X^{\perp W},$ (7)
where $(\pi_{1},\pi_{2})$ are the coefficients on $(W_{1},W_{2})$ in the linear projection of $X$ on $(1,W_{1},W_{2})$, and $X^{\perp W}$ is the intercept plus the projection residual. Note that $\pi_{1}$ is also the coefficient on $W_{1}$ in the linear projection of $X$ on $(1,W_{1})$, since we assumed $W_{1}$ and $W_{2}$ are uncorrelated.
Substituting equation (7) into the definition of $\delta$ (equation (2)) shows that $\delta$ can be written as a double ratio:
$\delta$ $=\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2})}\Bigg{/}\frac{\operatorname*{cov}(X,\gamma_{1,\text{long}}W_{1})}{\operatorname*{var}(\gamma_{1,\text{long}}W_{1})}=\frac{\pi_{2}}{\pi_{1}}\Big{/}\frac{\gamma_{2,\text{long}}}{\gamma_{1,\text{long}}}.$
The numerator measures the relative strength of two associations: treatment and the omitted variable ($\pi_{2}$), and treatment and the observed variable ($\pi_{1}$). Similarly, the denominator measures the relative strength of two associations: the outcome and the omitted variable ($\gamma_{2,\text{long}}$), and the outcome and the observed variable ($\gamma_{1,\text{long}}$). Hence $\delta$ compares the relative importance of unobservables and observables for treatment to their relative importance for outcomes.
This property that $\delta$ is a ratio of ratios leads to some counter-intuitive interpretations of the magnitude of $\delta$. For example, suppose $\gamma_{2,\text{long}}=c\cdot\gamma_{1,\text{long}}$ and $\pi_{2}=c\cdot\pi_{1}$ for $c=10$. Since both $W_{1}$ and $W_{2}$ are normalized to have unit variance, this value of $c$ indicates that $W_{2}$ has an association with $X$ (and $Y$) that is 10 times stronger than that of $W_{1}$ with $X$ (and $Y$). However, the implied value of $\delta$ is 1, which is traditionally interpreted as meaning that there is “equal selection” based on $W_{2}$ and on $W_{1}$. This value of $\delta$ is unchanged if $c=1/10$, in which case $W_{2}$’s association with $X$ and $Y$ is 10 times weaker than $W_{1}$’s association with those same variables. In fact, the value of $\delta$ is equal to 1 for all $c\neq 0$ in this example. As Cinelli and Hazlett (2020) point out, the interpretation of $\delta$ in some papers is at odds with its double ratio property. For example Oster (2019, page 192) states that $\delta=1$ implies that “the unobservable[s] and observables are equally related to the treatment.” Their paper recommends continuing to use $R_{\text{long}}^{2}$ as one of the sensitivity parameters, but that $\delta$ should be replaced with RXW2
W1
2
[…]
, where RXW2
W1
2
[…]
is the unknown $R$-squared in a regression of $X$ on $W_{2}$, after partialling out $W_{1}$.
5.2 Exogeneity of the Controls
All of the sensitivity analyses we discussed in sections 24 maintain assumption A4. That assumption requires all observed variables to be uncorrelated with all omitted variables. In many applications, this assumption would be viewed as strong—most observed variables are themselves correlated and it is unlikely that the omitted variables happen to be uncorrelated with all of the observed covariates, especially when many observed covariates are included in the analysis. Oster argues this assumption can be relaxed by redefining $W_{2}$ to be $W_{2}^{\perp W_{1}}=W_{2}-\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}W_{1}$, which is uncorrelated with $W_{1}$ by definition. However, as Diegert et al. (2022) argue, this redefinition of $W_{2}$ substantively changes the interpretation of $\delta$. The interpretation changes because the replacement of $W_{2}$ by $W_{2}^{\perp W_{1}}$ requires replacing the coefficients $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$ with $(\widetilde{\gamma}_{1},\gamma_{2,\text{long}})$ where $\widetilde{\gamma}_{1}=\gamma_{1,\text{long}}+\gamma_{2,\text{long}}^{\prime}\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}$. The redefined $\delta$ is thus not only a function of the associations between $(W_{1},W_{2})$ and $X$ and $Y$, as discussed above, but also of the unknown covariance between $W_{2}$ and $W_{1}$. Concretely, the redefined $\delta$ is
$\delta=\frac{\pi_{2}}{\pi_{1}}\Bigg{/}\frac{\gamma_{2,\text{long}}}{\gamma_{1,\text{long}}+\gamma_{2,\text{long}}\frac{\operatorname*{cov}(W_{2},W_{1})}{\operatorname*{var}(W_{1})}}.$
This approach implies that applied researchers have to either (a) assume unobserved confounders are uncorrelated with all controls, or (b) reason about a parameter that depends on this unknown covariance. This additional modification, on top of the double ratio discussion above, makes the interpretation of this redefined $\delta$ even more difficult to interpret.
This concern can be addressed by using an alternative parameterization of the omitted variable bias formula, such as those in Cinelli and Hazlett (2020) or Diegert et al. (2022). The parameterization in Cinelli and Hazlett (2020) continues to use $R_{\text{long}}^{2}$ but replaces $\delta$ with RXW2
W1
2
[…]
2
[…]
(see section 5.1 above). The parameterization in Diegert et al. (2022) replaces both $\delta$ and $R_{\text{long}}^{2}$ with $r_{X}^{2}\coloneqq\operatorname*{var}(\pi_{2}^{\prime}W_{2})/\operatorname*{var}(\pi_{1}^{\prime}W_{1})$ and $r_{Y}^{2}\coloneqq\operatorname*{var}(\gamma_{2,\text{long}}^{\prime}W_{2})/\operatorname*{var}(\gamma_{1,\text{long}}^{\prime}W_{1})$, which are measures of the relative importance of $W_{2}$ and $W_{1}$ in the selection and outcome equation, respectively.
5.3 The Role of $R_{\text{long}}^{2}$
All of the analysis above assumes the exact value of $R_{\text{long}}^{2}$ is known. Researchers should also assess how their conclusions depend on the choice of $R_{\text{long}}^{2}$. This is not common practice, however; see our survey analysis in appendix LABEL:sec:inPractice. Instead, researchers commonly focus on one of two choices, either $1.3\widehat{R}_{\text{med}}^{2}$ or 1. The choice $R_{\text{long}}^{2}=1$ is often thought to be the most conservative. However, Basu (2022, propositions 6–8) shows that this is not correct. Specifically, he shows that the explain away breakdown point $\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})$ is not necessarily decreasing in $R_{\text{long}}^{2}$. Consequently, according to this measure of robustness, smaller values of $R_{\text{long}}^{2}$ could lead to smaller values of the breakdown point, and hence less robust results. This fact can be overlooked if researchers only use a single value of $R_{\text{long}}^{2}$. For researchers who want to use the sensitivity parameters $\delta$ and $R_{\text{long}}^{2}$ we recommend plotting estimates of the identified sets ¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ for a range of $R_{\text{long}}^{2}$ values. Although we omit this analysis for brevity, researchers could also consider replacing the known $R_{\text{long}}^{2}$ assumption with an assumption that it is only known to be bounded, $R_{\text{long}}^{2}\in[R_{\text{med}}^{2},\overline{R}_{\text{long}}^{2}]$ for a known value of $\overline{R}_{\text{long}}^{2}$, and then reporting the identified set for $\beta_{\text{long}}$ as a function of $\overline{R}_{\text{long}}^{2}$ and the upper bound $\overline{\delta}$ on $\delta$.
5.4 Optimistic Bias Correction
Finally, De Luca et al. (2019, page 220) note that, when there are multiple elements of (4) (values of $\beta_{\text{long}}$ consistent with the data and assumptions), by default the Stata package psacalc only reports the element that is closest to $\beta_{\text{med}}$. Since the other elements are equally consistent with the data and assumptions, they note that there is no a priori reason to believe that the true magnitude of omitted variable bias happens to be the smallest one. This concern is also addressed by plotting the set ¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ rather than performing bias corrections.
6 Empirical Application: Social Capital and the Rise of the Nazi Party
Social capital is typically viewed as a good which leads to positive outcomes for countries. Satyanath et al. (2017, JPE) instead argued that “social capital can undermine and help to destroy a democratic system” (page 482) by studying the rise of the Nazi Party (the National Socialist German Worker’s Party [NSDAP]) in 1933. To do this, they analyzed a dataset of $n=229$ towns/cities in Germany. The outcome variable $Y$ is a measure of how many people in that city joined the Nazi party. The treatment variable $X$ is a measure of social capital, based on “association density,” the number of social clubs and associations per 1,000 city inhabitants. They consider three variations of this measure based on the type of association included: All associations, Civic only, and Military only. They also include a variety of socioeconomic and political covariates.
Their paper has a variety of analyses. We focus on their main results, which use a selection on observables identification strategy paired with linear models. That is, they estimate linear regressions of $Y$ on $X$, an intercept, and additional observed covariates. To assess the robustness of those results to omitted variables, they applied Oster (2019) in their subsection 4.D, with additional details in their appendix G. We re-examine that analysis in light of our theoretical results in section 3. See appendix LABEL:sec:extraEmpirical for extra details and results for this empirical application.
Breakdown Analysis
[…]
Reported Correct +$\infty$+ ∞ $10|\widehat{\beta}_{\text{med}}|$ $3|\widehat{\beta}_{\text{med}}|$ $2|\widehat{\beta}_{\text{med}}|$
$R_{\text{long}}^{2}$ (1) (2) (3) (4) (5) (6)
All $1$ $<0$ $-32.0$ 0.586 0.586 0.586 0.676
$1.3\widehat{R}_{\text{med}}^{2}$ $-331.5$ 0.953 0.954 1.476 2.829
Civic $1$ $<0$ 0.98 0.736 0.736 0.736 0.745
$1.3\widehat{R}_{\text{med}}^{2}$ 9.78 0.994 1.036 1.806 2.986
Military $1$ $<0$ 1.42 0.899 0.899 0.909 0.977
$1.3\widehat{R}_{\text{med}}^{2}$ 10.65 0.993 1.035 1.821 3.058
[…]
Table 6 shows the main breakdown results. This table replicates and extends the bottom half of Panel B in Table A.27 of appendix G in Satyanath et al. (2017). The three rows of Table 6 correspond to the three different versions of treatment. All of the cell entries are different kinds of breakdown points. First consider columns (1) and (2). These show estimates of Oster’s explain away breakdown point. We present estimates for two different choices of $R_{\text{long}}^{2}$, either 1 or Oster’s rule of thumb $1.3\widehat{R}_{\text{med}}^{2}$. Column (1) shows the values reported by Satyanath et al. (2017). As described in appendix LABEL:sec:commonMistake, these values are incorrect. Column (2) reports the correct values. If we judged robustness by comparing the magnitude of these values of the explain away breakdown point to 1, then all of the results would be deemed robust except one, civic associations with $R_{\text{long}}^{2}=1$, which gives an explain away breakdown point of 0.98.
However, as we have discussed in this paper, the explain away breakdown point can be very different from the sign change breakdown point. Columns (3)–(6) show estimated sign change breakdown points, for four increasingly restrictive choices of the constraint $M$ on the magnitude of OVB: +$\infty$+ ∞ (no constraint), $10|\widehat{\beta}_{\text{med}}|$, $3|\widehat{\beta}_{\text{med}}|$, and $2|\widehat{\beta}_{\text{med}}|$. First consider column (3), which does not impose any magnitude restriction. As implied by Theorem 2, we see that all of the point estimates are bounded above by 1.
As explained in section 3.2, we can obtain sign change breakdown points larger than 1 if we impose additional restrictions. Columns (4)–(6) show the estimated sign change breakdown points when we restrict the magnitude of OVB to be below $M$. For the relatively unrestrictive value of $M=10|\widehat{\beta}_{\text{med}}|$, this additional restriction does not change the sign change breakdown points much. However, it is enough to push two of the breakdown points slightly above 1. For the more restrictive values $M=3|\widehat{\beta}_{\text{med}}|$ and $M=2|\widehat{\beta}_{\text{med}}|$, the breakdown points are now all moderately larger than 1 for the rule of thumb choice $R_{\text{long}}^{2}=1.3\widehat{R}_{\text{med}}^{2}$. For $R_{\text{long}}^{2}=1$, however, the magnitude restrictions do not affect the values of the sign change breakdown points very much. Finally, notice that even with the strictest magnitude restriction, the sign change breakdown points (column 6) are substantially smaller than the explain away breakdown points (column 2).
Selection Bias Adjustments
{talltblr}
[…]
Baseline Estimate $\delta=0$ 0.170
Oster’s Prop 1 (our Prop LABEL:prop:OsterP1) $\delta=1$, LABEL:assump:OsterA2 0.532
[…]
Corollary 2, ^$\widehat{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1$ $(-\infty,-0.0855]\cup[0.0432,1.8947]$
Corollary 2, ^$\widehat{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1.05$ $(-\infty,-0.0798]\cup[0.0415,\infty)$
[…]
Satyanath et al. (2017) did not present selection bias adjusted estimates; as we document in appendix LABEL:sec:inPractice, it is rare for empirical researchers to present both breakdown analysis and bias adjustments. To illustrate the results of section 4, here we also use bias adjustments to analyze the sensitivity of the results in Satyanath et al. (2017).
Table 6 summarizes our bias adjustment results. For brevity, here we focus on just the “all associations” treatment variable and the specification in column (2) of appendix table LABEL:table:mainTable0, and we only consider $R_{\text{long}}^{2}=1$. Our findings are similar for the other treatment variables, specifications, and $R_{\text{long}}^{2}$ values. For simplicity we also do not present confidence sets for these point estimates. Panel A represents current empirical practice. The first row shows the baseline estimate. The second row shows the selection bias adjusted estimate obtained from Oster’s Proposition 1 (our Prop. LABEL:prop:OsterP1), which assumes $\delta=1$ and requires an additional assumption (LABEL:assump:OsterA2, discussed in appendix LABEL:sec:BiasCorrectionsAppendix). Based on this estimate, the researcher would conclude that selection bias only works in their favor—it has the same sign as their baseline estimate and indeed is even larger.
However, now consider the results in Panel B. If we maintain $\delta=1$ but drop LABEL:assump:OsterA2, then there are now two values of the coefficient that are consistent with the assumptions and the data, and one of them is negative. Hence even this minor change from current practice overturns the robustness finding based on Oster’s Proposition 1 alone. Now, a researcher may claim that the value $-0.0855$ is close to zero, and perhaps is not statistically significantly different from zero, in which case one may argue that Oster’s Proposition 2 with $\delta=1$ still weakly supports a positive effect. So next we consider perturbing $\delta$ a slight amount: If we set $\delta=0.99$ then the estimated identified set now includes a large negative value, $-18.66$. Likewise, if we set $\delta=1.01$ then the estimated identified set includes a large positive value, $15.64$. Thus we see that the magnitude of the bias adjustment is very sensitive to the specific choice of $\delta$. Finally, the last two rows of Panel B show the estimated identified sets when we replace the assumption that $\delta$ is known exactly with the assumption that we only have a bound $\bar{\delta}$ on its magnitude (A6). Here we see that for $\bar{\delta}=1$, almost every negative value is consistent with the data and the assumptions, and many positive values are as well, including some bigger and smaller than the baseline estimate. If we increase this bound slightly to $\bar{\delta}=1.05$ then arbitrarily large values of the coefficient are possible too.
Identified Set Plots
To better understand the estimates in tables 6 and 6, it is useful to plot the estimated identified set ^$\widehat{\mathcal{B}}_{I}(\delta,R_{\text{long}}^{2})$ as a function of $\delta$. Figure 1, which we discussed in the introduction, does this for the “all associations” specification from Satyanath et al. (2017) with $R_{\text{long}}^{2}=1$; this corresponds to the first row of Table 6. Figure 2 plots the corresponding estimated cumulative identified set as a function of $\bar{\delta}$, which assumes that $|\delta|\leq\bar{\delta}$, as well as its convex hull. These bounds provide a modified sensitivity analysis method that accurately represents the impact of omitted variables on the range of possible coefficient values. Our main recommendation is that empirical researchers present these set estimates. Note also that the magnitude constraint A5 can be imposed by simply superimposing horizontal lines at the largest and smallest feasible values on this plot.
7 Explaining Away versus Sign Changes in Practice: Two Meta-Analyses
In section 6 we illustrated the difference between explain away and sign change breakdown points in a single application. In this section, we study this difference in 58 different empirical applications by conducting two complementary meta-analyses. Overall, we show that this distinction between explaining away and sign changes leads to substantively different conclusions about robustness to omitted variables in the majority of empirical applications.
7.1 Data
The first meta-analysis is based on the papers we survey in appendix LABEL:sec:inPractice that were published in top five economics journals. We were able to replicate 12 of the 23 papers that implemented Oster’s methods; the other 11 papers either had private data or other hurdles that prevented a full replication. Among the 12 papers we replicated, we gathered all regressions where the authors implemented Oster’s methods, giving 194 total regressions. Within each paper, the robustness results for many of those regressions will be strongly correlated; for example, because they use the same outcome and treatment variables but with different controls. Hence we manually selected a subset of regressions to focus on, such as only including the longest specification per table. This gives our primary sample of 34 regressions from 12 papers.
The second meta-analysis is based on papers used in the meta-analysis of Oster (2019, sections 2 and 5). This dataset has 141 regressions from 55 different papers that were published in either a top 5 economics journal or the AEJ: Applied Economics between 2008 and 2013, and which reported results from coefficient stability analyses. See Oster (2019) for more details on the construction of this dataset. Since this dataset already only contains only a handful of regressions per paper, we did not further restrict the sample. This second dataset is useful both because it provides a larger sample of regressions and also because it the inclusion criteria are qualitatively different, since none of these papers implemented Oster’s methods because they preceded her paper.
7.2 Results
We present two sets of results. First, consider table 7.2. For each regression we estimated three breakdown points. First is the explain away breakdown point, also known as the absolute value of “Oster’s delta,” which represents current practice. Then we estimated two versions of the sign change breakdown point: One without any restriction on the magnitude of the OVB ($M=+\infty$) and the other which assumes the OVB is at most ten times the baseline estimate ($M=10\cdot|\widehat{\beta}_{\text{med}}|$). Table 7.2 shows percentiles of the distribution of these breakdown points across the various regressions for each of the two samples described above (panels A and B), and for the two most common choices of $R_{\text{long}}^{2}$ (Oster’s rule of thumb ($\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$), and 1).
[…]
$R^{2}_{\text{long}}$ Breakdown Point Type   10% 25% 50% 75% 90%
[…]
$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ Explain away 1.07 1.43 2.46 6.37 13.22
Sign change ($M=10|\widehat{\beta}_{\text{med}}|$) 0.78 1.03 1.21 1.47 1.66
Sign change ($M=+\infty$) 0.78 1.0 1.0 1.0 1.0
[…]
Sign change ($M=10|\widehat{\beta}_{\text{med}}|$) 0.21 0.22 0.88 1.17 1.37
Sign change ($M=+\infty$) 0.21 0.22 0.88 1.0 1.0
[…]
$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ Explain away 1.29 2.65 9.0 32.54 74.14
Sign change ($M=10|\widehat{\beta}_{\text{med}}|$) 0.95 1.15 2.21 8.53 31.52
Sign change ($M=+\infty$) 0.75 0.96 1.0 1.0 1.0
[…]
Sign change ($M=10|\widehat{\beta}_{\text{med}}|$) 0.88 1.13 1.79 3.48 5.24
Sign change ($M=+\infty$) 0.42 0.94 1.0 1.0 1.0
[…]
  • Note: Both samples only include results considered robust according to standard practice (the explain away breakdown point is weakly larger than 1), where we use the value of $R_{\text{long}}^{2}$ specified in the corresponding panel). Panel A: Top: $N=29$ regressions from 11 papers. Bottom: $N=11$ regressions from 6 papers. Panel B: Top: $N=111$ regressions from 47 papers. Bottom: $N=58$ regressions from 31 papers.
Here we further restrict the sample to all regressions that would be considered robust according to standard practice, meaning that the explain away breakdown point is larger than 1. By definition, the sign change breakdown point is always weakly smaller than the explain away breakdown point. Hence results that are not robust according to the explain away breakdown point will also be non-robust according to the sign change breakdown point. So this sample restriction allows us to answer the question: Among all results considered robust by standard practice, what is the impact of using sign change breakdown points to assess robustness?
First consider the sample in Panel A. As shown in Theorem 2, the sign change breakdown point without any magnitude restrictions is never larger than 1. This result can be seen in the last line of all panels. For the rule of thumb choice of $R_{\text{long}}^{2}$, most of these breakdown points exactly equal 1. For $R_{\text{long}}^{2}=1$, more than half of the regressions are no longer robust when using $\delta=1$ as the robustness threshold point. In practice, researchers almost always go beyond binary robust / not-robust conclusions and also report the magnitude of the explain away breakdown point, interpreting larger magnitudes as evidence of greater robustness. The second row of each sub-panel shows that even with a reasonably strong restriction on the magnitude of OVB, the sign change breakdown points are substantially smaller than the explain away breakdown points. For example, the median is cut in half. Near the top of the distribution, the 90th percentile shrinks 87% from 13.22 down to 1.66 for the rule of thumb $R_{\text{long}}^{2}$, and similarly for $R_{\text{long}}^{2}=1$. The results for Panel B are qualitatively similar: The sign change breakdown points are substantially smaller in magnitude than the explain away breakdown points. Overall, these results show that in most empirical applications, common conclusions like
“selection on unobservables would have to be at least 4.1 times stronger than selection on observables to explain away the relationship” (Squicciarini 2020, AER, page 20)
will often greatly exaggerate the magnitude of robustness, since they are based on explain away rather than sign change breakdown points (in this application, with $M=10|\widehat{\beta}_{\text{med}}|$ the sign change breakdown point is instead only 1.07; see table LABEL:table:squic).
[…]
$R^{2}_{\text{long}}$ $\delta_{\text{hypo}}$   10% 25% 50% 75% 90%
[…]
$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ $\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ 1.0 1.0 1.0 1.3 67.2
2 1.0 1.4 3.4 6.9 67.2
1.05 2.0 9.2 25.2 54.0 198.9
1 2.1 709.8 $\infty$ $\infty$ $\infty$
[…]
2 1.0 1.0 1.3 13.8 21.3
1.05 1.2 1.2 4.5 53.2 147.8
1 1.2 1.2 4.9 $\infty$ $\infty$
[…]
$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ $\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ 1.0 1.0 1.0 2.7 11.4
2 1.0 3.3 12.8 58.6 116.0
1.05 3.7 16.8 70.7 255.0 771.6
1 5.3 59.6 1978.6 $\infty$ $\infty$
[…]
2 1.0 2.1 10.8 45.8 211.8
1.05 1.5 16.3 50.5 192.3 760.0
1 1.6 32.1 216.4 $\infty$ $\infty$
[…]
  • Note: Both samples only include results considered robust according to standard practice (the explain away breakdown point is weakly larger than 1, where we use the value of $R_{\text{long}}^{2}$ specified in the corresponding panel). Panel A: Top: $N=29$ regressions from 11 papers. Bottom: $N=11$ regressions from 6 papers. Panel B: $N=111$ regressions from 47 papers. Bottom: $N=58$ regressions from 31 papers. The value $\infty$ means our optimizer to compute $\widehat{p}_{i}^{\text{bp}}$ stopped at the preset upper bound of 1 million, which can be interpreted as effectively infinite. For $\delta_{\text{hypo}}=2$ and 1.05, we actually use $\min\{2,\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})\}$ and $\min\{1.05,\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})\}$ to ensure that the distributions in each sub-panel are stochastically ranked from top to bottom.
Next consider table 7.2. Here we ask the question:
Consider a researcher who does not want to change their current practice at all, and continues to report the explain away breakdown point. What is the weakest possible restriction on the magnitude of OVB that is necessary to correctly interpret the explain away breakdown point as a sign change breakdown point?
To make magnitudes of OVB comparable across different regressions, we work with magnitudes relative to estimates of the baseline estimand $\beta_{\text{med}}$. Specifically, for the $i$th regression, let $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ be the largest magnitude of OVB (defined as a proportion of the baseline estimate) such that the sign change breakdown point is at least as large as $\delta_{\text{hypo}}$. Setting $\delta_{\text{hypo}}=\delta_{i}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ gives the answer to our question above—when can the explain away breakdown point be interpreted as a sign change breakdown point? [^4] Formally, we define $p_{i}^{\text{bp}}(\delta_{\text{hypo}})\coloneqq\sup\{p\geq 1:|\delta_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\beta_{\text{med},i}|)|\geq|\delta_{\text{hypo}}|\}$. Denote its estimator by $\widehat{p}_{i}^{\text{bp}}(\delta_{\text{hypo}})\coloneqq\sup\{p\geq 1:|\widehat{\delta}_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\widehat{\beta}_{\text{med},i}|)|\geq|\delta_{\text{hypo}}|\}$.
The first row of all four sub-panels in table 7.2 give the distribution of $\widehat{p}_{i}^{\text{bp}}(\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2}))$ across regressions $i$ in our meta-analysis samples, and for different choices of $R_{\text{long}}^{2}$. Here we see that, across both samples and both choices of $R_{\text{long}}^{2}$, more than 50% of regressions require that the sign of the parameter of interest $\beta_{\text{long}}$ must be known a priori if we want to interpret the explain away breakdown point as a sign change breakdown point. That follows since—recall from section 3.2—the choice $M=\beta_{\text{med}}$ (equivalently, $p=1$) implies that the magnitude restriction A5 is equivalent to $\beta_{\text{long}}\in[0,2\beta_{\text{med}}]$ (when the baseline estimand is positive; the negative case is analogous). If the sign of $\beta_{\text{long}}$ was known a priori, however, then there would be no need to perform a sensitivity analysis to assess whether omitted variables could flip the sign of one’s baseline results; they cannot by assumption.
For regressions above the 50th percentile, the implicit magnitude restrictions is typically larger than $p=1$, but is often still quite restrictive. For example, the 75th percentile in panel A is 1.3 under the rule of thumb $R_{\text{long}}^{2}$ choice. To interpret this, suppose $\widehat{\beta}_{\text{med}}=2$. Then $p=1.3$ corresponds to the assumption that $\beta_{\text{long}}\in[-0.6,4.6]$.
So far we have focused on the magnitude restriction required to interpret the explain away breakdown point as a sign change breakdown point. We can use $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ to answer several related questions as well. For example, what is the weakest possible restriction on the magnitude of OVB to ensure that the sign change breakdown point is at least 1? $p_{i}(\delta_{\text{hypo}})$ with $\delta_{\text{hypo}}=1$ answers this question. The bottom row of all sub-panels shows the distribution of our estimates of these magnitudes. Here we see that even for this very weak conclusion, at least 10% of regressions still require substantial magnitude restrictions. Most papers, however, require a fairly weak, or essentially no magnitude restriction in order to obtain a sign change breakdown point equal to or larger than 1. However, because of the asymptote at $\delta=1$, these $p_{i}^{\text{bp}}$ values may not be informative about values for $\delta$ cutoffs slightly above 1. So the middle rows of each sub-panel show estimates of $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ for values of $\delta_{\text{hypo}}$ close to, but slightly above 1. In these middle rows we see that obtaining sign change breakdown points which are meaningfully larger than 1 often requires substantial magnitude restrictions.
8 Conclusion
In this paper we distinguished between two kinds of breakdown points: sign change and explain away. Using theoretical results that apply regardless of the data generating process (section 3), four detailed empirical applications (section 6 and appendix LABEL:sec:extraEmpirical), and two meta-analyses covering 58 papers (section 7), we showed that the distinction between these two types of breakdown points matters in most empirical settings. In particular, explain away breakdown points often substantially overstate the robustness of results to omitted variables that can flip their sign. Given these results, we recommend that researchers plot the estimated identified set for the coefficient of interest under a variety of different assumptions about omitted variables; for example, see figure 2. We also recommend reporting sign change breakdown points as robustness summary statistics, such as in table 6. Both of these recommendations are easily implemented by using our companion Stata package regsensitivity.

### v4

@@
 (2005a, b, 2008).
-It was later substantially extended by Oster (2019) (hereafter Oster), an extremely influential paper with about 4500 Google Scholar citations as of December 2024.
+It was later substantially extended by Oster (2019) (hereafter Oster), an extremely influential paper with about 4900 Google Scholar citations as of June 2025.
 Finkelstein et al.
@@
 (2005a), Clarke (2009), Bellows and Miguel (2009), Hosman et al.
-(2010), Gonzalez and Miguel (2015), Krauth (2016), Cinelli and Hazlett (2020), and Diegert et al.
-(2022) for linear models and Rosenbaum and Rubin (1983), Rosenbaum (1995, 2002), Robins et al.
+(2010), Gonzalez and Miguel (2015), Krauth (2016), Cinelli and Hazlett (2020), Diegert et al.
+(2025a, b) for linear models and Rosenbaum and Rubin (1983), Rosenbaum (1995, 2002), Robins et al.
 (2000), Altonji, Elder, and Taber (2005a, 2008), Masten and Poirier (2018, 2020), and Altonji et al.
@@
 In fact, we show that any time the explain away breakdown point is large—suggesting that omitted variables may be unimportant—a much smaller value of her sensitivity parameter can actually change the sign of the parameter of interest.
-In particular, we prove that the sign change breakdown point is bounded above by $\delta=1$.
+In particular, we prove that the sign change breakdown point is bounded above by $1$.
 This is a substantial concern since 1 is often viewed as the cutoff for a robust result.
@@
 This explains the source of the discontinuity in the value of the omitted variable bias that we discussed earlier, and why for this sensitivity parameter it is easier to flip the coefficient’s sign than to obtain an exact zero.
-Concretely, in this example the baseline estimate is on the vertical intercept at $\delta=0$.
-This estimate is positive.
+Concretely, the baseline estimate is obtained at $\delta=0$.
+In this empirical example, the baseline estimate is positive.
 How robust is this estimate to the presence of omitted variables?
@@
 However, from the right plot, we see that we only need to go to $\delta=0.586$ to find a value of $\beta$ that is negative.
-This value $\delta=0.586$ is the sign change breakdown point.
-Oster (2019) does not discuss this breakdown point, its companion Stata package psacalc does not compute it, and none of the top five papers we survey in appendix LABEL:sec:inPractice report it.
+This value $|\delta|=0.586$ is the sign change breakdown point.
+Oster (2019) does not discuss this breakdown point, its companion Stata package psacalc does not compute it, and none of the papers we survey in Appendix LABEL:sec:inPractice report it.
 Yet, as we see here, the sign change breakdown point is substantially smaller than $|-32|$, the explain away breakdown point.
-Indeed, it is below $\delta=1$, the commonly used cutoff for a robust result.
+Indeed, its magnitude is below $1$, the commonly used cutoff for a robust result.
 Consequently, in this empirical example, focusing on the sign change breakdown point overturns the authors’ conclusion of robustness: Only a small amount of selection on unobservables relative to observables is necessary to flip the sign of the baseline estimate.
@@
 If the magnitude is restricted too much, then this assumption is essentially equivalent to assuming a priori that there is no omitted variable bias, in which case a sensitivity analysis would be unnecessary.
-But as soon as reasonably sized magnitudes are allowed, the concerns we discuss in sections 3 and 4 arise.
+But as soon as reasonably sized magnitudes are allowed, the concerns we discuss in sections 3 and 4 can arise.
 Refer to caption         Refer to caption
@@
 The graph of these values of $\beta$ is discontinuous in $\bar{\delta}$ at $\bar{\delta}=0.586$, which explains the difference in the sign change breakdown and explain away breakdown points.
-As we saw in figure 1, nearly all values of $\beta$, including arbitrarily large negative values, are consistent with an upper bound of $\bar{\delta}=1.05$, but a small range around 0 is not part of this set until $\bar{\delta}$ equals 32.
+As we saw in Figure 1, nearly all values of $\beta$, including arbitrarily large negative values, are consistent with an upper bound of $\bar{\delta}=1.05$, but a small range around 0 is excluded from the set, and the value 0 does not become part of this set until $\bar{\delta}$ equals 32.
 Researchers who want to focus only on the largest and smallest possible values of $\beta$ can instead present the convex hull, which is shown in the right plot of Figure 2.
@@
 We also illustrate the extreme sensitivity of bias adjusted estimates to perturbations of the sensitivity parameter.
-In appendix LABEL:sec:extraEmpirical we study three additional empirical applications in detail, all published in the American Economic Review.
+In Appendix LABEL:sec:extraEmpirical_applications we study three additional empirical applications in detail, all published in the American Economic Review.
 To see that these four applications are not outliers, in Section 7 we conduct two meta-analyses covering 58 different empirical applications published in the top five economics journals plus the AEJ: Applied Economics.
@@
 In particular, suppose researchers were to continue the current standard practice of reporting explain away breakdown points, but wanted to interpret them as sign change breakdown points.
-We show that in more than 50% of applications, this interpretation is only valid if researchers already know the sign of coefficient of interest.
+We show that in more than 50% of applications, this interpretation is only valid if researchers already know the sign of the coefficient of interest before performing the sensitivity analysis.
 But in that case, it would not be necessary to do a sensitivity analysis, since the result is known a priori.
@@
 Let $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$ denote the coefficients on $W\coloneqq(W_{1},W_{2})$.
-Suppose $\beta_{\text{long}}$ is the parameter of interest. In particular, we can write the long regression as
+It is well known that linear regression coefficients can be given causal interpretations under various identifying assumptions (e.g., Angrist and Pischke 2009).
+In this paper, we take it as given that $\beta_{\text{long}}$ is the parameter of interest.
+This focus on $\beta_{\text{long}}$ is often because we have in fact made assumptions that make it a causal parameter, but our technical results do not rely on those motivating assumptions.
+So, taking interest in $\beta_{\text{long}}$ as given, write the long regression as
 $Y=\beta_{\text{long}}X+\gamma_{1,\text{long}}^{\prime}W_{1}+\gamma_{2,\text{long}}^{\prime}W_{2}+Y^{\perp X,W}$ (1)
@@
 Several papers propose alternative sensitivity parameters, including Cinelli and Hazlett (2020) and Diegert et al.
-(2022).
+(2025a).
 In this paper we instead use the same sensitivity parameter $\delta$ as Oster (2019), since our focus here is to study the implications of using different kinds of breakdown points.
@@
 Denote these baseline controls by $W_{0}$.
-To incorporate these covariates, replace $(X,W_{1},W_{2})$ with $(X^{\perp W_{0}},W_{1}^{\perp W_{0}},W_{2}^{\perp W_{0}})$ throughout the entire analysis.
+To incorporate these covariates, replace $(X,W_{1},W_{2})$ with $(X^{\perp W_{0}},W_{1}^{\perp W_{0}},W_{2}^{\perp W_{0}})$, the variables after residualizing with respect to $W_{0}$, throughout the entire analysis.
 See Section 3.4 of Diegert et al.
-(2022) for a justification of this approach and further discussion.
+(2025a) for a justification of this approach and further discussion.
 For simplicity, we omit $W_{0}$ throughout our theoretical analysis.
@@
 Appendix A.1 in Oster (2019) briefly discusses one approach to relaxing A4, which we discuss in Section 5.
-Also see Appendix A of Diegert et al.
-(2022) for additional details.
+Also see Appendix C of Diegert et al.
+(2025b) for additional details.
 Diegert et al.
-(2022) also develop an alternative approach to sensitivity analysis that does not rely on A4.
+(2025a) develop an alternative approach to sensitivity analysis that does not rely on A4.
 In this paper we maintain A4 throughout because our main focus is to study different notions of breakdown.
 For simplicity, from here on we assume $W_{2}$ is a scalar.
-All of our results can be generalized to vector $W_{2}$ using the arguments in Appendix B of Diegert et al.
-(2022).
+All of our results can be generalized to vector $W_{2}$ by working with the scalar index $\gamma_{2,\text{long}}^{\prime}W_{2}$ everywhere (as in the proof of Theorem 2 in Diegert et al.
+2025a, for example).
 To characterize the possible values for $\beta_{\text{long}}$, it is useful to consider the regression of $X$ on $(1,W_{1})$.
@@
 Consequently, the absolute value of this term equals $\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})$.
+The proof of Theorem 1 and all other formal results can be found in Appendix LABEL:sec:proofs.
 Theorem 1 is very similar to Oster’s Proposition 3, with only minor differences.
@@
 Oster’s delta answers the following question:
-  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $\delta$ that is required for the data to be consistent with $\beta_{\text{long}}=0$?
+  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $|\delta|$ that is required for the data to be consistent with $\beta_{\text{long}}=0$?
 We call this the explain away breakdown point. Researchers may also be interested in answering the following question:
-  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $\delta$ such that $\beta_{\text{long}}$ has a different sign from $\beta_{\text{med}}$?
+  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $|\delta|$ such that $\beta_{\text{long}}$ has a different sign from $\beta_{\text{med}}$?
 We call this the sign change breakdown point.
@@
 Theorem 2 shows that, in Oster’s sensitivity analysis, the sign change breakdown point can never be larger than one.
-In practice, researchers typically consider the value $\delta=1$ to be the cutoff for determining the robustness of a result.
-That is, they interpret values of $\delta$ greater than 1 as robust and values less than 1 as sensitive.
-Based on this criterion, Theorem 2 therefore shows that the sign of the parameter of interest is never robust (assuming we require $\delta$ strictly larger than 1 to conclude that the result is robust).
+In practice, researchers typically consider the value $|\delta|=1$ to be the cutoff for determining the robustness of a result.
+That is, they interpret values of $|\delta|$ greater than 1 as robust and values less than 1 as sensitive.
+Based on this criterion, Theorem 2 therefore shows that the sign of the parameter of interest is never robust (assuming we require $|\delta|$ strictly larger than 1 to conclude that the result is robust).
 In contrast, the explain away breakdown point $\delta^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ is often substantially larger than one.
@@
     $\delta$’s measure of selection on unobservables is the coefficient from regressing $X$ on $\gamma_{2,\text{long}}W_{2}$:
-    $\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2})}.$
+    $\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2})}.$ (6)
     $\gamma_{2,\text{long}}$, as one of the long regression coefficients, will also be large because of near multi-collinearity.
 Consequently, this measure of selection on unobservables will be close to zero since regression coefficients scale down when the covariate is scaled up.
+More formally, note that $\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})$ and $\operatorname*{var}(\gamma_{2,\text{long}}W_{2})$ both diverge to infinity as we approach perfect multi-collinearity since $|\gamma_{2,\text{long}}|\rightarrow\infty$ in this scenario.
+However their ratio (the regression coefficient in (6)) converges to zero since its numerator is linear in $\gamma_{2,\text{long}}$ while its denominator is quadratic in $\gamma_{2,\text{long}}$.
   3. 3.
@@
 This is exactly what the alternative methods of Cinelli and Hazlett (2020) and Diegert et al.
-(2022) do.
+(2025a) do.
 Both papers include sensitivity parameters that explicitly bound the joint explanatory power of the observed and unobserved covariates in the first stage.
+The sensitivity analyses proposed by both papers then lead to sign change and explain away breakdown points that are equal.
 3.2 Incorporating Magnitude Restrictions on the Omitted Variable Bias
@@
 Suppose our estimate of $\beta_{\text{med}}$ is 0.1.
-Then $M=1$ would say that the true coefficient on schooling—as obtained after adjusting for the unobserved confounders—is at most 1.1.
-This value is far outside the range of most estimates (e.g., Card 2001 table II) and so in this application researchers would likely view A5 with $M=1$ as a plausible assumption.
+Then $M=1$ would say that the true coefficient on schooling—as obtained after adjusting for the unobserved confounders—lies between $-0.9$ and 1.1.
+This interval includes most estimates (e.g., Card 2001 Table II) so in this application researchers would likely view A5 with $M=1$ as a plausible assumption.
 In general, a simple approach to selecting $M$ is to let it be a multiple of $\beta_{\text{med}}$. That is, let $M=p\cdot|\beta_{\text{med}}|$ for $p\geq 0$. For this choice, A5 is equivalent to
-$\beta_{\text{long}}\in\big{[}\beta_{\text{med}}-p|\beta_{\text{med}}|,\beta_{\text{med}}+p|\beta_{\text{med}}|\big{]}.$
-Every element of this set has the same sign if and only if $p<1$.
+$\beta_{\text{long}}\in\big{[}\beta_{\text{med}}-p\cdot|\beta_{\text{med}}|,\beta_{\text{med}}+p\cdot|\beta_{\text{med}}|\big{]}.$
+All elements of this set have the same sign if $p<1$.
 For this reason, since we are interested in robustness to sign changes, $p$ should be chosen to be larger than 1.
@@
 Recall that A2 specifies that $\delta\in\mathbb{R}$ is known.
-Then the identified set for $\beta_{\text{long}}$ is
-$\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M)\coloneqq\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\cap[\beta_{\text{med}}-M,\beta_{\text{med}}+M].$
+Denote the identified set for $\beta_{\text{long}}$ under these assumptions by $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M)$.
+Then
+$\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M)=\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\cap[\beta_{\text{med}}-M,\beta_{\text{med}}+M].$
 Let $\delta^{\text{bp},\text{sign}}(R_{\text{long}}^{2},M)$ denote the corresponding sign change breakdown point.
@@
 Suppose $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known.
-Then the identified set for $\beta_{\text{long}}$ is
-¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})\coloneqq\bigcup_{|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2}).$
-Suppose A5 also holds. Then the identified set for $\beta_{\text{long}}$ is
-¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2},M)\coloneqq\bigcup_{|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M).$
+Denote the identified set for $\beta_{\text{long}}$ under these assumptions by ¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$.
+Then
+¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})=\bigcup_{\delta:|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2}).$
+Suppose A5 also holds. Denote the identified set for $\beta_{\text{long}}$ under this as well as the earlier assumptions by ¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2},M)$. Then
+¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2},M)=\bigcup_{\delta:|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M).$
 This corollary simply notes that we take the union of the identified sets from equation (5) over all allowed values of the sensitivity parameter $\delta$ with magnitude less than $\bar{\delta}$ to obtain the identified set under A6.
@@
 Hence $\delta$ compares the relative importance of unobservables and observables for treatment to their relative importance for outcomes.
-This property that $\delta$ is a ratio of ratios leads to some counter-intuitive interpretations of the magnitude of $\delta$.
-For example, suppose $\gamma_{2,\text{long}}=c\cdot\gamma_{1,\text{long}}$ and $\pi_{2}=c\cdot\pi_{1}$ for $c=10$.
-Since both $W_{1}$ and $W_{2}$ are normalized to have unit variance, this value of $c$ indicates that $W_{2}$ has an association with $X$ (and $Y$) that is 10 times stronger than that of $W_{1}$ with $X$ (and $Y$).
-However, the implied value of $\delta$ is 1, which is traditionally interpreted as meaning that there is “equal selection” based on $W_{2}$ and on $W_{1}$.
-This value of $\delta$ is unchanged if $c=1/10$, in which case $W_{2}$’s association with $X$ and $Y$ is 10 times weaker than $W_{1}$’s association with those same variables.
-In fact, the value of $\delta$ is equal to 1 for all $c\neq 0$ in this example.
-As Cinelli and Hazlett (2020) point out, the interpretation of $\delta$ in some papers is at odds with its double ratio property.
-For example Oster (2019, page 192) states that $\delta=1$ implies that “the unobservable[s] and observables are equally related to the treatment.” Their paper recommends continuing to use $R_{\text{long}}^{2}$ as one of the sensitivity parameters, but that $\delta$ should be replaced with RXW2
-W1
-2
 […]
-[…]
-[…]
-, where RXW2
-W1
-2
-[…]
-[…]
-[…]
-is the unknown $R$-squared in a regression of $X$ on $W_{2}$, after partialling out $W_{1}$.
 5.2 Exogeneity of the Controls
@@
 In many applications, this assumption would be viewed as strong—most observed variables are themselves correlated and it is unlikely that the omitted variables happen to be uncorrelated with all of the observed covariates, especially when many observed covariates are included in the analysis.
-Oster argues this assumption can be relaxed by redefining $W_{2}$ to be $W_{2}^{\perp W_{1}}=W_{2}-\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}W_{1}$, which is uncorrelated with $W_{1}$ by definition.
+Oster argues this assumption can be relaxed by redefining $W_{2}$ to be $W_{2}^{\perp W_{1}}:=W_{2}-\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}W_{1}$, which is uncorrelated with $W_{1}$ by definition.
 However, as Diegert et al.
-(2022) argue, this redefinition of $W_{2}$ substantively changes the interpretation of $\delta$.
-The interpretation changes because the replacement of $W_{2}$ by $W_{2}^{\perp W_{1}}$ requires replacing the coefficients $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$ with $(\widetilde{\gamma}_{1},\gamma_{2,\text{long}})$ where $\widetilde{\gamma}_{1}=\gamma_{1,\text{long}}+\gamma_{2,\text{long}}^{\prime}\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}$.
+(2025b) argue, this redefinition of $W_{2}$ substantively changes the interpretation of $\delta$.
+The interpretation changes because the replacement of $W_{2}$ by $W_{2}^{\perp W_{1}}$ requires replacing the coefficients $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$ with $(\widetilde{\gamma}_{1},\gamma_{2,\text{long}})$ where $\widetilde{\gamma}_{1}:=\gamma_{1,\text{long}}+\operatorname*{var}(W_{1})^{-1}\operatorname*{cov}(W_{1},W_{2})\gamma_{2,\text{long}}$.
 The redefined $\delta$ is thus not only a function of the associations between $(W_{1},W_{2})$ and $X$ and $Y$, as discussed above, but also of the unknown covariance between $W_{2}$ and $W_{1}$.
-Concretely, the redefined $\delta$ is
-$\delta=\frac{\pi_{2}}{\pi_{1}}\Bigg{/}\frac{\gamma_{2,\text{long}}}{\gamma_{1,\text{long}}+\gamma_{2,\text{long}}\frac{\operatorname*{cov}(W_{2},W_{1})}{\operatorname*{var}(W_{1})}}.$
+Concretely, for scalar $W_{1}$ and $W_{2}$, the redefined $\delta$ is
+$\widetilde{\delta}$ $\coloneqq\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2}^{\perp W_{1}})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2}^{\perp W_{1}})}\Bigg{/}\frac{\operatorname*{cov}(X,\widetilde{\gamma}_{1}W_{1})}{\operatorname*{var}(\widetilde{\gamma}_{1}W_{1})}$
+$=\frac{\pi_{2}}{\pi_{1}+\pi_{2}\frac{\operatorname*{cov}(W_{2},W_{1})}{\operatorname*{var}(W_{1})}}\Bigg{/}\frac{\gamma_{2,\text{long}}}{\gamma_{1,\text{long}}+\gamma_{2,\text{long}}\frac{\operatorname*{cov}(W_{2},W_{1})}{\operatorname*{var}(W_{1})}}.$
 This approach implies that applied researchers have to either (a) assume unobserved confounders are uncorrelated with all controls, or (b) reason about a parameter that depends on this unknown covariance.
 This additional modification, on top of the double ratio discussion above, makes the interpretation of this redefined $\delta$ even more difficult to interpret.
-This concern can be addressed by using an alternative parameterization of the omitted variable bias formula, such as those in Cinelli and Hazlett (2020) or Diegert et al.
-(2022).
-The parameterization in Cinelli and Hazlett (2020) continues to use $R_{\text{long}}^{2}$ but replaces $\delta$ with RXW2
-W1
-2
 […]
-[…]
-2
-[…]
-(see section 5.1 above).
-The parameterization in Diegert et al.
-(2022) replaces both $\delta$ and $R_{\text{long}}^{2}$ with $r_{X}^{2}\coloneqq\operatorname*{var}(\pi_{2}^{\prime}W_{2})/\operatorname*{var}(\pi_{1}^{\prime}W_{1})$ and $r_{Y}^{2}\coloneqq\operatorname*{var}(\gamma_{2,\text{long}}^{\prime}W_{2})/\operatorname*{var}(\gamma_{1,\text{long}}^{\prime}W_{1})$, which are measures of the relative importance of $W_{2}$ and $W_{1}$ in the selection and outcome equation, respectively.
 5.3 The Role of $R_{\text{long}}^{2}$
@@
 We re-examine that analysis in light of our theoretical results in Section 3.
-See appendix LABEL:sec:extraEmpirical for extra details and results for this empirical application.
+See Appendix LABEL:sec:extraEmpirical_details for extra details and results for this empirical application.
 Breakdown Analysis
@@
 (2017).
-The three rows of Table 6 correspond to the three different versions of treatment.
+The three panels of Table 6 correspond to the three different versions of treatment.
 All of the cell entries are different kinds of breakdown points.
@@
 […]
-Corollary 2, ^$\widehat{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1$ $(-\infty,-0.0855]\cup[0.0432,1.8947]$
-Corollary 2, ^$\widehat{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1.05$ $(-\infty,-0.0798]\cup[0.0415,\infty)$
+Corollary 2, ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1$ $(-\infty,-0.0855]\cup[0.0432,1.8947]$
+Corollary 2, ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1.05$ $(-\infty,-0.0798]\cup[0.0415,\infty)$
 […]
@@
 If we increase this bound slightly to $\bar{\delta}=1.05$ then arbitrarily large values of the coefficient are possible too.
-Identified Set Plots
+Identified Set Plots and Main Empirical Recommendations
 To better understand the estimates in tables 6 and 6, it is useful to plot the estimated identified set ^$\widehat{\mathcal{B}}_{I}(\delta,R_{\text{long}}^{2})$ as a function of $\delta$.
@@
 Note also that the magnitude constraint A5 can be imposed by simply superimposing horizontal lines at the largest and smallest feasible values on this plot.
+Researchers may also want to summarize the information shown in plots like Figure 2 of the cumulative identified set.
+We recommend two ways in which researchers can do this, both of which modify and extend current practice as documented in Appendix LABEL:sec:inPractice.
+  1. 1.
+    First, in lieu of computing bias adjusted estimates based on equation (7), researchers can estimate the cumulative identified set ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\overline{\delta},R_{\text{long}}^{2})$ for a reference value of $\overline{\delta}$, such as $\overline{\delta}=1$.
+Then they can select an empirically plausible value of $M$, the largest allowed value of the omitted variable bias.
+Then remove any elements from ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\overline{\delta},R_{\text{long}}^{2})$ that lay outside the range $[\widehat{\beta}_{\text{med}}-M,\widehat{\beta}_{\text{med}}+M]$, because these are deemed a priori implausible.
+The set of all remaining elements is then the estimated set of long regression coefficient values consistent with $|\delta|\leq\overline{\delta}$, the choice of $R_{\text{long}}^{2}$, and the restriction that the magnitude of OVB is no larger than $M$.
+This set can be interpreted as a set of bias corrected coefficients.
+For example, consider the specification in the second row of Table 6, with $M=2|\widehat{\beta}_{\text{med}}|=2\times 0.17=0.34$ (column (6)).
+In this case, ¯^$\widehat{\overline{\mathcal{B}}}_{I}(1,1.3\widehat{R}_{\text{med}}^{2})\cap[-0.17,0.51]=\big{(}(-\infty,-0.84]\cup[0.17,0.19]\big{)}\cap[-0.17,0.51]=[0.17,0.19]$.
+Thus researchers can state:
+    “Assuming the omitted variable bias is no larger than $0.34$, and that selection on unobservables is no larger than selection on observables, the bias corrected set of coefficient estimates is $[0.17,0.19]$.”
+  2. 2.
+    Second, in lieu of computing explain away breakdown points, researchers can estimate the sign change breakdown point for an empirically plausible value of $M$.
+For example, consider again the specification in the second row of table 6, with $M=2|\widehat{\beta}_{\text{med}}|=0.34$ (column (6)).
+Here the estimated sign change breakdown point is 2.83.
+Thus researchers can state:
+    “Assuming the omitted variable bias is no larger than 0.34, we estimate that selection on unobservables would need to be at least 2.83 times as large as selection on observables to change the sign of our baseline result.”
+Whenever summary statements like these are made, we recommend that researchers discuss their choice of $M$, report their choice of $R_{\text{long}}^{2}$ (and consider exploring additional choices, see Section 5.3), and also still present identified set plots like in Figure 2, if only in an appendix.
 7 Explaining Away versus Sign Changes in Practice: Two Meta-Analyses
@@
 Since this dataset already only contains only a handful of regressions per paper, we did not further restrict the sample.
-This second dataset is useful both because it provides a larger sample of regressions and also because it the inclusion criteria are qualitatively different, since none of these papers implemented Oster’s methods because they preceded her paper.
+This second dataset is useful both because it provides a larger sample of regressions and also because the inclusion criteria are qualitatively different, since none of these papers implemented Oster’s methods because they preceded her paper.
 7.2 Results
@@
 For the rule of thumb choice of $R_{\text{long}}^{2}$, most of these breakdown points exactly equal 1.
-For $R_{\text{long}}^{2}=1$, more than half of the regressions are no longer robust when using $\delta=1$ as the robustness threshold point.
+For $R_{\text{long}}^{2}=1$, more than half of the regressions are no longer robust when using $1$ as the robustness threshold point.
 In practice, researchers almost always go beyond binary robust / not-robust conclusions and also report the magnitude of the explain away breakdown point, interpreting larger magnitudes as evidence of greater robustness.
@@
 Overall, these results show that in most empirical applications, common conclusions like
-“selection on unobservables would have to be at least 4.1 times stronger than selection on observables to explain away the relationship” (Squicciarini 2020, AER, page 20)
+“selection on unobservables would have to be at least 4.1 times stronger than selection on observables to explain away the relationship” (Squicciarini 2020, AER, page 20) [Note: This computation uses $R_{\text{long}}^{2}=1$]
 will often greatly exaggerate the magnitude of robustness, since they are based on explain away rather than sign change breakdown points (in this application, with $M=10|\widehat{\beta}_{\text{med}}|$ the sign change breakdown point is instead only 1.07; see Table LABEL:table:squic).
@@
 […]
-$R^{2}_{\text{long}}$ $\delta_{\text{hypo}}$   10% 25% 50% 75% 90%
+$R^{2}_{\text{long}}$ 10% 25% 50% 75% 90%
 […]
-$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ $\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ 1.0 1.0 1.0 1.3 67.2
-2 1.0 1.4 3.4 6.9 67.2
-1.05 2.0 9.2 25.2 54.0 198.9
-1 2.1 709.8 $\infty$ $\infty$ $\infty$
+$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ 1.0 1.0 1.0 1.3 67.2
+1 1.0 1.0 1.0 13.8 21.3
 […]
-2 1.0 1.0 1.3 13.8 21.3
-1.05 1.2 1.2 4.5 53.2 147.8
-1 1.2 1.2 4.9 $\infty$ $\infty$
-[…]
-$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ $\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ 1.0 1.0 1.0 2.7 11.4
-2 1.0 3.3 12.8 58.6 116.0
-1.05 3.7 16.8 70.7 255.0 771.6
-1 5.3 59.6 1978.6 $\infty$ $\infty$
-[…]
-2 1.0 2.1 10.8 45.8 211.8
-1.05 1.5 16.3 50.5 192.3 760.0
-1 1.6 32.1 216.4 $\infty$ $\infty$
+$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ 1.0 1.0 1.0 2.7 11.4
+1 1.0 1.0 1.0 3.9 61.6
 […]
   • Note: Both samples only include results considered robust according to standard practice (the explain away breakdown point is weakly larger than 1, where we use the value of $R_{\text{long}}^{2}$ specified in the corresponding panel).
-Panel A: Top: $N=29$ regressions from 11 papers.
-Bottom: $N=11$ regressions from 6 papers.
-Panel B: $N=111$ regressions from 47 papers.
-Bottom: $N=58$ regressions from 31 papers.
-The value $\infty$ means our optimizer to compute $\widehat{p}_{i}^{\text{bp}}$ stopped at the preset upper bound of 1 million, which can be interpreted as effectively infinite.
-For $\delta_{\text{hypo}}=2$ and 1.05, we actually use $\min\{2,\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})\}$ and $\min\{1.05,\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})\}$ to ensure that the distributions in each sub-panel are stochastically ranked from top to bottom.
+Panel A: First row: $N=29$ regressions from 11 papers.
+Second row: $N=11$ regressions from 6 papers.
+Panel B: First row: $N=111$ regressions from 47 papers.
+Second row: $N=58$ regressions from 31 papers.
 Next consider Table 7.2. Here we ask the question:
@@
 What is the weakest possible restriction on the magnitude of OVB that is necessary to correctly interpret the explain away breakdown point as a sign change breakdown point?
-To make magnitudes of OVB comparable across different regressions, we work with magnitudes relative to estimates of the baseline estimand $\beta_{\text{med}}$.
-Specifically, for the $i$th regression, let $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ be the largest magnitude of OVB (defined as a proportion of the baseline estimate) such that the sign change breakdown point is at least as large as $\delta_{\text{hypo}}$.
-Setting $\delta_{\text{hypo}}=\delta_{i}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ gives the answer to our question above—when can the explain away breakdown point be interpreted as a sign change breakdown point?
-[^4] Formally, we define $p_{i}^{\text{bp}}(\delta_{\text{hypo}})\coloneqq\sup\{p\geq 1:|\delta_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\beta_{\text{med},i}|)|\geq|\delta_{\text{hypo}}|\}$.
-Denote its estimator by $\widehat{p}_{i}^{\text{bp}}(\delta_{\text{hypo}})\coloneqq\sup\{p\geq 1:|\widehat{\delta}_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\widehat{\beta}_{\text{med},i}|)|\geq|\delta_{\text{hypo}}|\}$.
-The first row of all four sub-panels in table 7.2 give the distribution of $\widehat{p}_{i}^{\text{bp}}(\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2}))$ across regressions $i$ in our meta-analysis samples, and for different choices of $R_{\text{long}}^{2}$.
+To make magnitudes of OVB comparable across different regressions, we work with magnitudes relative to the baseline estimates $\widehat{\beta}_{\text{med}}$.
+Specifically, for the $i$th regression, let $\widehat{p}_{i}^{\text{bp}}$ be the largest magnitude of OVB (defined as a proportion of the baseline estimate) such that the estimated sign change breakdown point is at least as large as the estimated explain away breakdown point in this regression.
+[^4] Formally, we define $\widehat{p}_{i}^{\text{bp}}\coloneqq\sup\{p\geq 1:|\widehat{\delta}_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\widehat{\beta}_{\text{med},i}|)|\geq|\widehat{\delta}^{\text{bp},\text{explain away}}_{i}(R_{\text{long}}^{2})|\}$.
+The two panels in Table 7.2 give the distribution of $\widehat{p}_{i}^{\text{bp}}$ across regressions $i$ in our meta-analysis samples for different choices of $R_{\text{long}}^{2}$.
 Here we see that, across both samples and both choices of $R_{\text{long}}^{2}$, more than 50% of regressions require that the sign of the parameter of interest $\beta_{\text{long}}$ must be known a priori if we want to interpret the explain away breakdown point as a sign change breakdown point.
@@
 If the sign of $\beta_{\text{long}}$ was known a priori, however, then there would be no need to perform a sensitivity analysis to assess whether omitted variables could flip the sign of one’s baseline results; they cannot by assumption.
-For regressions above the 50th percentile, the implicit magnitude restrictions is typically larger than $p=1$, but is often still quite restrictive.
+For regressions above the 50th percentile, the implicit magnitude restriction is typically larger than $p=1$, but is often still quite restrictive.
 For example, the 75th percentile in Panel A is 1.3 under the rule of thumb $R_{\text{long}}^{2}$ choice.
@@
 Then $p=1.3$ corresponds to the assumption that $\beta_{\text{long}}\in[-0.6,4.6]$.
-So far we have focused on the magnitude restriction required to interpret the explain away breakdown point as a sign change breakdown point.
-We can use $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ to answer several related questions as well.
-For example, what is the weakest possible restriction on the magnitude of OVB to ensure that the sign change breakdown point is at least 1?
-$p_{i}(\delta_{\text{hypo}})$ with $\delta_{\text{hypo}}=1$ answers this question.
-The bottom row of all sub-panels shows the distribution of our estimates of these magnitudes.
-Here we see that even for this very weak conclusion, at least 10% of regressions still require substantial magnitude restrictions.
-Most papers, however, require a fairly weak, or essentially no magnitude restriction in order to obtain a sign change breakdown point equal to or larger than 1.
-However, because of the asymptote at $\delta=1$, these $p_{i}^{\text{bp}}$ values may not be informative about values for $\delta$ cutoffs slightly above 1.
-So the middle rows of each sub-panel show estimates of $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ for values of $\delta_{\text{hypo}}$ close to, but slightly above 1.
-In these middle rows we see that obtaining sign change breakdown points which are meaningfully larger than 1 often requires substantial magnitude restrictions.
+Additional related results for these meta-analyses are in Appendix LABEL:sec:extra_meta-analyses.
 8 Conclusion
 In this paper we distinguished between two kinds of breakdown points: sign change and explain away.
-Using theoretical results that apply regardless of the data generating process (section 3), four detailed empirical applications (section 6 and appendix LABEL:sec:extraEmpirical), and two meta-analyses covering 58 papers (section 7), we showed that the distinction between these two types of breakdown points matters in most empirical settings.
+Using theoretical results that apply regardless of the data generating process (Section 3), four detailed empirical applications (Section 6 and Appendix LABEL:sec:extraEmpirical_applications), and two meta-analyses covering 58 papers (Section 7), we showed that the distinction between these two types of breakdown points matters in most empirical settings.
 In particular, explain away breakdown points often substantially overstate the robustness of results to omitted variables that can flip their sign.